# Augmented Lagrangian Method based adjoint space framework for sparse reconstruction of acoustic source with boundary measurements

### v1

@@ 3.1 Semismooth Newton-based augmented Lagrangian framework in the dual space
The final step size $t_{l}:=\beta^{t}$ obtained from the line search in (3.28) dictates the update of $y^{k+1}$ as follows:
@@
Once some stopping criterion for inner Newton iteration is satisfied, e.g., for some $l=l_{0}$, we set $y^{k+1}=y^{l_{0}}$ and obtain $z^{k+1}$ with $z^{k+1}=\mathcal{M}^{k}(y^{k+1})$. After finishing the inner semismooth Newton iterations, we then update $\lambda^{k+1}$ as in (3.14b) for the outer ALM updates. Once the stopping criterion of the outer ALM iterations is satisfied with $(y^{K},z^{K},\lambda^{K})$, the solution $\mu$ corresponding to the primal problem can be approximated by
@@
which has an explicit solution. Since by the primal-dual optimality conditions of (3.10) at saddle-points $(y^{*},\mu^{*})$, we have
@@
Algorithm 1 Boundary measurements based dual Augmented Lagrangian method with semismooth Newton solver for (2.8) (ALM-bd-SSN)
 Given linear operator $\mathcal{V}_{b}$ and $u_{b}$, regularization parameters $(\alpha_{0},\alpha)$, and parameter settings for ALM $\sigma_{0}$, $c_{0}$, $\beta$, $\rho$, and initial vectors $(y_{0},z_{0},\lambda_{0})$,
@@ 3.2 Convergence of the SSN-based ALM on the dual space
Now we introduce some basic definitions and properties of multivalued mapping from convex analysis [20, 37]. Let $F:X\rightrightarrows Y$ be a multivalued mapping. The graph of $F$ is defined as the set
@@
The inverse of $F$, i.e., $F^{-1}:Y\rightrightarrows X$ is defined as the multivalued mapping whose graph is $\{(y,x)|(x,y)\in\text{gph}F\}$. The distance $x$ from the set $C\subset X$ is defined by
@@
A mapping $F:X\rightrightarrows Y$ is called metrically subregular at $\bar{x}$ for $\bar{y}$ if $(\bar{x},\bar{y})\in\text{gph}F$ and there exists modulus $\kappa\geq 0$ along with a neighborhoods $U$ of $\bar{x}$ and $V$ of $\bar{y}$ such that
@@
Now, let us turn to the primal problem (3.1), which is the dual problem (3.10). For the $L^{1}$ norm, we see
@@
It is well-known that $l$ is a convex-concave function on $(u,p,\lambda)$. Define the maximal monotone operator $T_{l}$ by
@@
For the dual problem (3.10), assuming the KKT system has at least one solution, then $T_{l}$ is metrically subregular at $(y^{*},z^{*},\lambda^{*})^{T}$ for the origin. Similarly, assuming $(\partial\mathcal{P})^{-1}(0)\neq\emptyset$ as $\mathcal{P}$ in (3.1), $\partial\mathcal{P}$ are metrically subregular at $\lambda^{*}$ for the origin.
@@
It is known that the $\|\cdot\|_{1}$ (3.32) is a polyhedral convex function. We can see that $\partial h^{*}$ is a polyhedral function of $y$ and $\partial p^{*}$ is a polyhedral function of $z$. We thus conclude that the monotone operator $\mathcal{A}$ is polyhedral. Besides, the operator $\mathcal{B}$ is a maximal monotone and linear operator. Thus $T_{l}$ is a polyhedral mapping [40]. By the corollary in [40], we see $T_{l}$ is metrically subregular at $(y^{*},z^{*},\lambda^{*})^{T}$ for the origin.
Let us now turn to the metric subregularity of $\partial\mathcal{P}$, which is a dual function of the dual problem (3.10). Supposing $(\partial\mathcal{P})^{-1}(0)\neq\emptyset$, since
@@
Here, we follow the standard stopping criterion for the inexact augmented Lagrangian method originated [41, 42].
@@
Compared to (3.6), we conclude that optimal solutions $\mu^{*}$ of (3.6) and $\lambda^{*}$ of (3.40) have the relation $\mu^{*}=-\lambda^{*}$. The convergence of $\mu^{k}$ is dominated by $\lambda^{k}$. We can also obtain the convergence rate of $\mu^{k}$ from the convergence rate of $y^{k}$ by (3.30). Since $(I+\frac{\alpha}{\alpha_{0}}\partial\|.\|_{1})^{-1}$ is a firmly nonexpansive operator, we have
@@
We denote $\mathcal{X}^{P}$ as the solution sets of the problem (3.1). With these preparations, we have the following global convergence and local convergence rate for the ALM method (3.14), under the mild condition that the KKT system of (3.33) has at least one solution, as in Theorem 2.
@@
Since discrete $L^{2}(\partial\Omega)$ is a finite-dimensional reflexive space and the dual function (3.10) is lower semicontinuous, proper, and strongly convex due to the strong convexity of $h^{*}$, it is coercive. Thus, the existence of the solution can be guaranteed [28, Theorem 4.25]. Furthermore, since $\operatorname{dom}\mathcal{D}=L^{2}(\partial\Omega)$, by Fenchel-Rockafellar theory [28, Chapter 4.3], the solution to the dual problem (3.10) is not empty and $\min_{\mu}\mathcal{P}(\mu)=\max_{y}-\mathcal{D}(y)$. By [42, Theorem 4] (or [41, Theorem 1] where the augmented Lagrangian method (3.14) essentially comes from proximal point method applying to its dual problem, i.e., the primal problem (3.1)), with criterion (A), we get the boundedness of $\{\lambda^{k}\}_{k}$. The uniqueness of $(y^{*},z^{*})$ follows from the strong convexity of $h^{*}$ on $y$ and the $z^{*}=-\mathcal{V}_{b}^{T}y^{*}$, which is one of the KKT conditions. The boundedness of $(y^{k},z^{k})$ and convergence of $(y^{k},z^{k},\lambda^{k})$ then follows by [42, Theorem 4].
@@ 4 Semismooth Newton Method and First-Order Primal-dual Method
In this section, we will introduce a semismooth Newton method and a first-order primal-dual method for the inverse source scattering problem. Let us first discuss the semismooth Newton method. For the primal formulation (3.1), if choosing
@@
where $B:=\mathcal{V}_{b}^{T}\mathcal{V}_{b}+\alpha_{0}I$. The the primal formulation can be written as $\min_{\mu}f(B\mu)+g(\mu)$. By the Fenchel–Rockafellar duality, the predual problem is $\min_{y}f^{*}(y)+g^{*}(-B^{T}y)$, which is
@@
Although there is $B^{-1}$ in the above functional (4.3), with direct calculation by the first-order optimality condition of the functional (4.3), no $B^{-1}$ is involved in the following nonlinear equation, fortunately,
@@
The formulation $\max(0,B^{T}y-\alpha)$ or $\min(0,B^{T}y+\alpha)$ is understood in point-wise sense. For their subgradients, we introduce $\mathcal{A}_{k}^{+}=\{x\in\Omega:B^{T}y^{k}(x)\geq\alpha\}$, $\mathcal{A}_{k}^{-}=\{x\in\Omega:B^{T}y^{k}(x)\leq-\alpha\}$, and $\mathcal{A}_{k}=\mathcal{A}_{k}^{+}\cup\mathcal{A}_{k}^{-}$ along with $\chi_{\mathcal{A}^{+}}=\text{Diag}([\chi_{\mathcal{A}^{+}}]_{1},[\chi_{\mathcal{A}^{+}}]_{2},\ldots,)$, $\chi_{\mathcal{A}^{-}}=\text{Diag}(\chi_{\mathcal{A}^{1}}]_{1},[\chi_{\mathcal{A}^{1}}]_{2},\ldots,)$ and $\chi_{\mathcal{A}}=\text{Diag}([\chi_{\mathcal{A}}]_{1},[\chi_{\mathcal{A}}]_{2},\ldots,)$ depend on $y$ and their diagonal elements are defined by
@@
We let $\gamma\rightarrow\infty$ with path-following strategy [25] (or [28, Section 9.1]). Finally, we can recover $\mu$ with $y$ by (3.5)
@@
Algorithm 2 Semismooth Newton method (shortened as SSN)
@@ 4.1 First-order primal-dual method
Finally, let us turn to the first-order methods. The Chambolle-Pock’s first-order primal-dual algorithm [14] is employed and studied for the inverse medium scattering problem in [12, 13]. Here, for comparison, we also presented the first-order primal-dual algorithm. With
@@
We now can employ the first-order primal-dual algorithm [14, Algorithm 1] to solve (2.8) with saddle-point formulation (4.10), as the following Algorithm 3.
Algorithm 3 First-order primal-dual algorithm (PDA)
@@ 5 Numerical Experiments
The volume potential $\mathcal{V}_{k}$ and $(I-\mathcal{V}_{k}q)^{-1}$ can be computed by the Fourier-based collocation method by periodization of the Lippmann-Schwinger volume potential [31]. We adapted the code developed in [12, 13] where the discretization of $\mathcal{T}$ is given, and we refer to [12, 13] for more details.
All algorithms are conducted with MATLAB R2024b on a Ubuntu 24.04 workstation with dual Intel (R) Xeon (R) E5-2697AV4 CPUs, 2.60 GHz, and 288 GB memory. The parameters of the compared algorithms are as follows
  • ALM: ALM-bd-SSN as in Algorithm 1. We choose $\sigma_{0}=1$ and $\sigma_{k+1}=6\sigma_{k}$. For the Armijo line search, we choose $\beta=0.3$ and $c=10^{-4}$ as in (3.28).
@@
The Gaussian noise is added to get noisy boundary measurement $u_{b}^{\delta}$ for the simulated data $u$, which is essentially as the same in [12, (73)]
@@
where $N_{\text{re}}$ and $N_{\text{im}}$ are two real matrices sampled from standard normal distribution. We measure the reconstruction error by the following relative error
@@
To avoid “inverse crime”, we employ the code of [12, 13]. The strategy is that the discretization, especially the grid size of generating the synthetic data of the direct problem, is different from the discretization for reconstructing the unknown sources [12] (see also [47, Chapter 2.3.6]).
For the homogeneous medium, we choose the velocity $c\equiv 1$ both in $\mathbb{R}^{2}$ and $\mathbb{R}^{3}$. For the inhomogeneous medium, we employ the inhomogeneous medium from [26]
@@
Figures 2 and 3 present the reconstructions of different point sources with %1 relative Gaussian noise error for homogeneous and inhomogeneous medium correspondingly. Table 1 collects the running time and relative errors of three different algorithms. For our experiments, we did not see big differences in running time between the homogeneous and the inhomogeneous medium. We thus only present the running time for the homogeneous case.
Figures 4 and 5 present the reconstructions of different strip sources with %0.1 relative Gaussian noise error for homogeneous and inhomogeneous medium correspondingly. Table 2 collects corresponding running time only for the homogeneous cases (still because the running time for the inhomogeneous cases is nearly the same as the homogeneous cases), relative errors for differential algorithms regarding homogeneous or inhomogeneous medium.
Figures 6 and 7 present the reconstructions of different 3-dimensional sources with %1 relative Gaussian noise error for homogeneous and inhomogeneous medium, respectively. Table 3 collects the running time and relative errors of ALM and SSN. It can be seen that the ALM can be 10 times faster than SSN. We did not compare with PDA since we found PDA is too slow. We computed the solution with $\lambda^{k}$ since $\mu^{*}=-\lambda^{*}$ by (3.40) and numerically we found that $\|\mu^{K_{\text{max}}}+\lambda^{K_{\text{max}}}\|$ is generally less than $10^{-5}$ as $K_{\text{max}}$} being the maximal outer iteration number and $\mu^{K_{\text{max}}}$ computed by (3.30).
Figures 8 and 9 present the reconstructions of different 3-dimensional sources with %0.01 relative Gaussian noise error for homogeneous and inhomogeneous medium, correspondingly. Table 4 collects relative errors of ALM and SSN. Since the running time to similar to the %1 relative Gaussian noise error case, we omit the running time in Table 4.
Figure 2: Reconstruction of sparse sources with multiple peaks in homogeneous media with $k=6$. The figures in the leftmost column are the original acoustic sources, including one, four, six, and eight peaks. The images in the second, the third from the left, and the rightmost columns are reconstructed results of ALM, SSN, and PDA, respectively. The images in the first, second, third, and fourth rows are the sources with one, four, six, and eight peaks, respectively.
@@
Table 1: Comparisons of the running time and relative errors for ALM, SSN, and PDA algorithms with noise level 1%. This table corresponds to Figure 2 and Figure 3. The running times, i.e., “Times (s)” are only presented for the homogeneous cases as in Figure 2. The “N-error” and “N-error (in)” are relative errors for the corresponding homogeneous and inhomogeneous cases. The notations “one”, “four”, “six”, and “eight” corresponds to the sources with one, four, six, and eight peaks and are the same as in Figures 2 and 3. (ALM and SSN with $\alpha$ = 9e-4, $\alpha_{0}$= 1e-7, and PDA with $\alpha$ = 9e-5, $\alpha_{0}$= 1e-12).
@@
Table 2: Comparisons of the running time and relative errors for ALM, SSN, and PDA algorithms with noise level 0.1%, and $\alpha$ = 2e-6, together with $\alpha_{0}$= 2e-8. This table corresponds to Figure 2 and Figure 3. The running times, i.e., “Times (s)” are only presented for the homogeneous cases as in Figure 4. The “N-error” and “N-error (in)” are relative errors for the corresponding homogeneous and inhomogeneous cases. The notations “skew” and “diag” represent the “skew diagonal” and “diagonal” strips in Figures 4 and 5.
@@
Figure 6: Reconstruction of 3-dimensional sparse sources in homogeneous media with $k=6$, noise level 1%, and $\alpha$ = 5e-7 together with $\alpha_{0}$= 2e-8. The solution is computed by $\lambda^{K_{\text{max}}}$. The figures in the leftmost column are the original acoustic sources. The images in the middle and the rightmost columns are reconstructed results of ALM and SSN, respectively. The images in the first, second, third, and fourth rows are the sources with the two balls, the right up tripod, the left down tripod, and the two tripods, respectively.
@@
Table 3: Experimental results for ALM and SSN methods under different scenarios and treatments, noise level 1%, and $\alpha$ = 5e-7 together with $\alpha_{0}$= 2e-8. Here “N-Error” is the relative error for both homogeneous and inhomogeneous cases, corresponding to “homo” and “inhomo” in the table. The notations “two balls”, “two tripods”, “right up”, and “left down” represent the acoustic sources in the first to the fourth rows in Figure 6 or 7, respectively.
@@
Figure 8: Reconstruction of 3-dimensional sparse sources in homogeneous media with $k=6$, noise level 0.01%, and $\alpha$ = 1e-7 together with $\alpha_{0}$= 1e-9. Reconstruction of 3-dimensional sparse sources in homogeneous media with $k=6$ and noise level 1%, and $\alpha$ = 5e-7 together with $\alpha_{0}$= 2e-8. The figures in the leftmost column are the original acoustic sources. The images in the middle and the rightmost columns are reconstructed results of ALM and SSN, respectively. The images in the first, second, and third rows are the sources with the two balls, the right up tripod, and the left down tripod, respectively.
@@
Table 4: Relative reconstruction errors for ALM and SSN methods with noise level 0.01%, and $\alpha$ = 1e-7 together with $\alpha_{0}$= 1e-9. The “N-error” and “N-error (in)” are relative errors for the corresponding homogeneous and inhomogeneous cases. The notations “two balls”, “right up”, and “left down” represent the acoustic sources in the first to the third rows in Figure 8 or 9, respectively. The running times are nearly the same to the 1% noise level cases as in Table and we omit here.

### v3

@@ 3.1 Semismooth Newton-based augmented Lagrangian framework in the dual space
The final step size $t_{l}:=\beta^{t}$ obtained from the line search in (3.28) is used to update of $y^{k+1}$ as follows:
@@
Once a stopping criterion for inner Newton iteration is satisfied, e.g., for some $l=l_{0}$, we set $y^{k+1}=y^{l_{0}}$ and obtain $z^{k+1}$ with $z^{k+1}=\mathcal{M}^{k}(y^{k+1})$. After finishing the inner semismooth Newton iterations, we then update $\lambda^{k+1}$ as in (3.14b) for the outer ALM updates. Once the stopping criterion of the outer ALM iterations is satisfied with $(y^{K},z^{K},\lambda^{K})$, the solution $\mu$ corresponding to the primal problem can be approximated by
@@
which has an explicit solution. Since, by the primal-dual optimality conditions of (3.10) at saddle-points $(y^{*},\mu^{*})$, we have
@@
Algorithm 1 Boundary Measurements-Based Dual Augmented Lagrangian Method with Semismooth Newton Solver for (2.8) (ALM-bd-SSN)
 Given linear operator $\mathcal{V}_{b}$ and $u_{b}$, regularization parameters $(\alpha_{0},\alpha)$, parameter settings for ALM $\sigma_{0}$, $c_{0}$, parameters for line search $\beta$, $c$, and initial vectors $(y_{0},z_{0},\lambda_{0})$, maximum outer ALM iteration number $K_{\text{max}}$, maximum SSN iteration number $L$.
@@ 3.2 Convergence of the SSN-based ALM on the dual space
Henceforth, we assume all the spaces including $L^{2}(\partial\Omega)$, $L^{2}(\Omega)$, all the variables including $y$, $\mu$, $z$, and $\lambda$, and all the linear operators including $\mathcal{V}_{b}$ and $\mathcal{V}_{b}^{T}$ are in finite-dimensional settings. Now we introduce some basic definitions and properties of multivalued mappings from convex analysis [20, 37]. Let $F:X\rightrightarrows Y$ be a multivalued mapping. The graph of $F$ is defined as the set
@@
The inverse of $F$, i.e., $F^{-1}:Y\rightrightarrows X$ is defined as the multivalued mapping whose graph is $\{(y,x)|(x,y)\in\text{gph}F\}$. The distance from $x$ to the set $C\subset X$ is defined by
@@
A mapping $F:X\rightrightarrows Y$ is called metrically subregular at $\bar{x}$ for $\bar{y}$ if $(\bar{x},\bar{y})\in\text{gph}F$ and there exists a modulus $\kappa\geq 0$ along with a neighborhood $U$ of $\bar{x}$ and $V$ of $\bar{y}$ such that
@@
Now, let us turn to the primal problem (3.1) and its dual problem (3.10). For the $L^{1}$ norm, we see
@@
It is well known that $l$ is a convex-concave function on $(y,z,\lambda)$. Define the maximal monotone operator $T_{l}$ by
@@
For the dual problem (3.10), assuming the KKT system has at least one solution, then $T_{l}$ is metrically subregular at $(y^{*},z^{*},\lambda^{*})^{T}$ for the origin. Similarly, assuming $(\partial\mathcal{P})^{-1}(0)\neq\emptyset$ as $\mathcal{P}$ in (3.1), $\partial\mathcal{P}$ is metrically subregular at $\lambda^{*}$ for the origin.
@@
It is known that the $\|\cdot\|_{1}$ (3.32) is a polyhedral convex function. We can see that $\partial h^{*}$ is a polyhedral function of $y$ and $\partial p^{*}$ is a polyhedral function of $z$. We thus conclude that the monotone operator $\mathcal{A}$ is polyhedral. Besides, the operator $\mathcal{B}$ is maximal monotone and linear. Thus $T_{l}$ is a polyhedral mapping [40]. By the corollary in [40], we see $T_{l}$ is metrically subregular at $(y^{*},z^{*},\lambda^{*})^{T}$ for the origin.
Let us now turn to the metric subregularity of $\partial\mathcal{P}$, the subdifferential of the dual problem (3.10). Assuming $(\partial\mathcal{P})^{-1}(0)\neq\emptyset$, since
@@
Here, we follow the standard stopping criteria for the inexact augmented Lagrangian method originated [41, 42].
@@
Compared to (3.6), we conclude that optimal solutions $\mu^{*}$ of (3.6) and $\lambda^{*}$ of (3.40) have the relation $\mu^{*}=-\lambda^{*}$. The convergence of $\mu^{k}$ is determined by $\lambda^{k}$. We can also obtain the convergence rate of $\mu^{k}$ from the convergence rate of $y^{k}$ by (3.30). Since $(I+\frac{\alpha}{\alpha_{0}}\partial\|.\|_{1})^{-1}$ is a firmly nonexpansive operator, we have
@@
We denote $\mathcal{X}^{P}$ as the solution set of the problem (3.1). With these preparations, we have the following global convergence and local convergence rate for the ALM method (3.14), under the mild condition that the KKT system of (3.33) has at least one solution, as in Theorem 2.
@@
Since discrete $L^{2}(\partial\Omega)$ is a finite-dimensional reflexive space and the dual function (3.10) is lower semicontinuous, proper, and strongly convex due to the strong convexity of $h^{*}$, it is coercive. Thus, the existence of the solution can be guaranteed [28, Theorem 4.25]. Furthermore, since $\operatorname{dom}\mathcal{D}=L^{2}(\partial\Omega)$, by Fenchel-Rockafellar theory [28, Chapter 4.3], the solution to the dual problem (3.10) is not empty and $\min_{\mu}\mathcal{P}(\mu)=\max_{y}-\mathcal{D}(y)$. By [42, Theorem 4] (or [41, Theorem 1], where the augmented Lagrangian method (3.14) essentially comes from proximal point method applying to its dual problem, i.e., the primal problem (3.1)), with criterion (A), we obtain the boundedness of $\{\lambda^{k}\}_{k}$. The uniqueness of $(y^{*},z^{*})$ follows from the strong convexity of $h^{*}$ on $y$ and the $z^{*}=-\mathcal{V}_{b}^{T}y^{*}$, which is one of the KKT conditions. The boundedness of $(y^{k},z^{k})$ and convergence of $(y^{k},z^{k},\lambda^{k})$ then follow by [42, Theorem 4].
@@ 4 Semismooth Newton method and first-order primal-dual method
In this section, we will introduce a semismooth Newton method and a first-order primal-dual method for the inverse source scattering problem. Let us first discuss the semismooth Newton method. For the primal formulation (3.1), by choosing
@@
where $B:=\mathcal{V}_{b}^{T}\mathcal{V}_{b}+\alpha_{0}I$. The primal formulation can be written as $\min_{\mu}f(B\mu)+g(\mu)$. By the Fenchel–Rockafellar duality, the predual problem is $\min_{y}f^{*}(y)+g^{*}(-B^{T}y)$, which is
@@
Although $B^{-1}$ appears in the functional (4.3), direct calculation of the first-order optimality condition shows that no $B^{-1}$ is involved in the following nonlinear equation, fortunately,
@@
The formulation $\max(0,B^{T}y-\alpha)$ or $\min(0,B^{T}y+\alpha)$ is understood componentwise. For their subgradients, we introduce $\mathcal{A}_{k}^{+}=\{x\in\Omega:B^{T}y^{k}(x)\geq\alpha\}$, $\mathcal{A}_{k}^{-}=\{x\in\Omega:B^{T}y^{k}(x)\leq-\alpha\}$, and $\mathcal{A}_{k}=\mathcal{A}_{k}^{+}\cup\mathcal{A}_{k}^{-}$ along with $\chi_{\mathcal{A}^{+}}=\text{Diag}([\chi_{\mathcal{A}^{+}}]_{1},[\chi_{\mathcal{A}^{+}}]_{2},\ldots,)$, $\chi_{\mathcal{A}^{-}}=\text{Diag}(\chi_{\mathcal{A}^{1}}]_{1},[\chi_{\mathcal{A}^{1}}]_{2},\ldots,)$ and $\chi_{\mathcal{A}}=\text{Diag}([\chi_{\mathcal{A}}]_{1},[\chi_{\mathcal{A}}]_{2},\ldots,)$, which depend on $y$ and their diagonal elements are defined by
@@
We let $\gamma\rightarrow\infty$ using a path-following strategy [25] (or [28, Section 9.1]). Finally, we can recover $\mu$ with $y$ by (3.5)
@@
Algorithm 2 Semismooth Newton Method (abbreviated as SSN)
@@ 4.1 First-order primal-dual method
Finally, let us turn to the first-order methods. The Chambolle-Pock first-order primal-dual algorithm [14] is employed and studied for the inverse medium scattering problem in [12, 13]. Here, for comparison, we also present the first-order primal-dual algorithm. With
@@
We now can employ the first-order primal-dual algorithm [14, Algorithm 1] to solve (2.8) in its saddle-point formulation (4.10), as in the following Algorithm 3.
Algorithm 3 First-Order Primal-Dual Algorithm (abbreviated as PDA)
@@ 5 Numerical experiments
The volume potential $\mathcal{V}_{k}$ and $(I-\mathcal{V}_{k}q)^{-1}$ can be computed using a Fourier-based collocation method via periodization of the Lippmann-Schwinger volume potential [31]. We adapted the code developed in [12, 13] where the discretization of $\mathcal{T}$ is given, and we refer to [12, 13] for more details.
All algorithms are implemented in MATLAB R2024b on an Ubuntu 24.04 workstation with dual Intel Xeon E5-2697 v4 CPUs (2.60 GHz) and 288 GB of memory. The parameters of the compared algorithms are as follows
  • ALM: ALM-bd-SSN as in Algorithm 1. We choose $\sigma_{0}=1$ and $\sigma_{k+1}=6\sigma_{k}$ with $c_{0}=6$. For the Armijo line search, we choose $\beta=0.3$ and $c=10^{-4}$ as in (3.28).
@@
Gaussian noise is added to the simulated data $u$ to obtain noisy boundary measurements $u_{b}^{\delta}$, which is essentially the same as in [12, (73)]
@@
where $N_{\text{re}}$ and $N_{\text{im}}$ are two real vectors sampled from standard normal distribution. We measure the reconstruction error by the following relative error
@@
To avoid “inverse crime”, we use the implementation from [12, 13]. The strategy is to use different discretizations (especially grid sizes) for generating the synthetic data in the direct problem and for reconstructing the unknown sources [12] (see also [47, Chapter 2.3.6]).
For the homogeneous medium, we choose the velocity $c\equiv 1$ both in $\mathbb{R}^{2}$ and $\mathbb{R}^{3}$. For the inhomogeneous medium, we employ the one from [26]
@@
Figures 2 and 3 show reconstructions of point sources with %1 relative Gaussian noise for homogeneous and inhomogeneous media, respectively. Table 1 collects the running time and relative errors of three different algorithms. For our experiments, we observed no significant differences in running time between the homogeneous and inhomogeneous cases. We thus only present the running time for the homogeneous case.
Figures 4 and 5 present the reconstructions of different strip sources with %0.1 relative Gaussian noise for homogeneous and inhomogeneous media, respectively. Table 2 collects the corresponding running times only for the homogeneous cases (since the running times for the inhomogeneous cases are nearly the same), and relative errors for the different algorithms in both homogeneous and inhomogeneous media.
Figures 6 and 7 present the reconstructions of different 3-dimensional sources with %1 relative Gaussian noise for homogeneous and inhomogeneous media, respectively. Table 3 collects the running time and relative errors of ALM and SSN. It can be seen that ALM can be 10 times faster than SSN. We did not compare with PDA, as we found it too slow. We computed the solution using $\lambda^{k}$ since $\mu^{*}=-\lambda^{*}$ by (3.40), and numerically we found that $\|\mu^{K_{\text{max}}}+\lambda^{K_{\text{max}}}\|$ is generally less than $10^{-5}$. Here $K_{\text{max}}$} is the maximal outer iteration number and $\mu^{K_{\text{max}}}$ is computed by (3.30).
Figures 8 and 9 present the reconstructions of different 3-dimensional sources with %0.01 relative Gaussian noise for homogeneous and inhomogeneous media, respectively. Table 4 collects relative errors of ALM and SSN. Since the running times are similar to those in the %1 relative Gaussian noise case, we omit the running time in Table 4.
Figure 2: Reconstruction of sparse sources with multiple peaks in homogeneous media with $k=6$. The images in the leftmost column are the original acoustic sources, including one, four, six, and eight peaks. The images in the second, the third from the left, and the rightmost columns are reconstructed results of ALM, SSN, and PDA, respectively. The images in the first, second, third, and fourth rows are the sources with one, four, six, and eight peaks, respectively.
@@
Table 1: Comparisons of the running time and relative errors for ALM, SSN, and PDA algorithms with noise level 1%. This table corresponds to Figures 2 and 3. The running times, i.e., “Times (s)” are only presented for the homogeneous cases as in Figure 2. The “N-error” and “N-error (in)” are relative errors for the corresponding homogeneous and inhomogeneous cases. The notations “one”, “four”, “six”, and “eight” correspond to the sources with one, four, six, and eight peaks and are the same as in Figures 2 and 3. (ALM and SSN with $\alpha$ = 9e-4, $\alpha_{0}$= 1e-7, and PDA with $\alpha$ = 9e-5, $\alpha_{0}$= 1e-12).
@@
Table 2: Comparisons of the running time and relative errors for ALM, SSN, and PDA algorithms with noise level 0.1%, and $\alpha$ = 2e-6, together with $\alpha_{0}$= 2e-8. This table corresponds to Figures 4 and 5. The running times, i.e., “Times (s)” are only presented for the homogeneous cases as in Figure 4. The “N-error” and “N-error (in)” are relative errors for the corresponding homogeneous and inhomogeneous cases. The notations “skew” and “diag” represent the “skew diagonal” and “diagonal” strips in Figures 4 and 5.
@@
Figure 6: Reconstruction of 3-dimensional sparse sources in homogeneous media with $k=6$, noise level 1%, and $\alpha$ = 5e-7 together with $\alpha_{0}$= 2e-8. The solution is computed by $\lambda^{K_{\text{max}}}$. The figures in the leftmost column are the original acoustic sources. The images in the middle and the rightmost columns are reconstructed results of ALM and SSN, respectively. The images in the first, second, third, and fourth rows are the sources with the two balls, the right-up tripod, the left-down tripod, and the two tripods, respectively.
@@
Table 3: Experimental results for ALM and SSN methods under different scenarios and treatments, noise level 1%, and $\alpha$ = 5e-7 together with $\alpha_{0}$= 2e-8. Here “N-Error” is the relative error for both homogeneous and inhomogeneous cases, corresponding to “homo” and “inhomo” in the table. The notations “two balls”, “two tripods”, “right up”, and “left down” represent the acoustic sources in the first to the fourth rows in Figures 6 and 7, respectively.
@@
Figure 8: Reconstruction of 3-dimensional sparse sources in homogeneous media with $k=6$, noise level 0.01%, and $\alpha$ = 1e-7 together with $\alpha_{0}$= 1e-9. The figures in the leftmost column are the original acoustic sources. The images in the middle and the rightmost columns are reconstructed results of ALM and SSN, respectively. The images in the first, second, and third rows are the sources with the two balls, the right-up tripod, and the left-down tripod, respectively.
@@
Table 4: Relative reconstruction errors for ALM and SSN methods with noise level 0.01%, and $\alpha$ = 1e-7 together with $\alpha_{0}$= 1e-9. The “N-error” and “N-error (in)” are relative errors for the corresponding homogeneous and inhomogeneous cases. The notations “two balls”, “right up”, and “left down” represent the acoustic sources in the first to the third rows in Figures 8 and 9, respectively. The running times are nearly the same as the 1% noise level cases as in Table 3, and we omit them here.